# Energy-Efficient Transformer Inference: Optimization Strategies for Time Series Classification

### v3

@@ Discussion
Inference Time and Energy Efficiency
@@
Energy-Accuracy Trade-off Analysis
@@
Conclusion
Our study systematically investigates energy-efficient optimization strategies for transformer-based time series classification. By integrating structured pruning and quantization, we achieve significant improvements in inference speed and energy consumption while preserving robust classification performance. Notably, the proposed framework adapts to varying dataset characteristics, demonstrating that even models with higher computational complexity can be effectively compressed for deployment in resource-constrained environments. Our results underscore the need to tailor optimization strategies to specific application domains and offer valuable insights for developing scalable, sustainable deep learning solutions in edge computing scenarios.

### v4

@@ 6.1 Impact on Classification Accuracy
6.2 Inference Time and Energy Efficiency
@@
6.3 Energy-Accuracy Trade-off Analysis
@@
7 Conclusion
Our study systematically investigates energy-efficient optimization strategies for transformer-based time series classification. By integrating structured pruning and quantization, significant improvements are achieved in inference speed and energy consumption while preserving robust classification performance. Notably, the proposed framework adapts to varying dataset characteristics, demonstrating that even models with higher computational complexity can be effectively compressed for deployment in resource-constrained environments. Our results underscore the need to tailor optimization strategies to specific application domains and offer valuable insights for developing scalable, sustainable deep learning solutions in edge computing scenarios.